OpenStudy (anonymous):

Is there a horizontal asymptote for the function f(x) = 2x^3 + 2 over (x-3)

OpenStudy (anonymous):

NO ! Because : \[\lim_{x\to+\infty}f(x)=+\infty\\\lim_{x\to-\infty}f(x)=-\infty\]

OpenStudy (anonymous):

The VA is at x = 3 and there is a parabolic asymptote at 2 x^2+6 x+18

OpenStudy (anonymous):

@SA5UK3 The question "Is there a horizontal asymptote ...?"

OpenStudy (anonymous):

I didn't think there was one. Thank you!!